OpenStudy (anonymous):

An owner can lease her building for $120,000 per year for three years. The explicit cost of maintaining the building is $40,000, and the implicit cost is $55,000. All revenues are received, and costs borne, at the end of each year. If the interest rate is 5 percent, determine the present value of the stream of: A. Accounting Profits B. Economic Profits

OpenStudy (anonymous):

A) Accounting profits dont take implicit costs into account, only "real" or quantifiable costs. Thus the present value of a 120,000 lease at 5% for three years with explicit costs of $40,000 maintenance is: PV = [ FV/(1+r)^n ] - (Explicit Cost) PV = 120000/(1.05^3) - (40000*3) B) same thing but add implicit costs ... PV = 120000/(1.05^3) - (40000*3) - (55000*3)